# Estimating Causal Effects for Binary Outcomes Using Per-Decision Inverse Probability Weighting

### v2

@@ Appendix C Proof of Identifiability Result
Therefore, to prove (24) it suffices to show that for any $1\leq u\leq\Delta$,
@@
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$ (25)
@@
So right hand side of (25) equals
$E\left\{\prod_{j=t+1}^{t+u-1}\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$ (26)
Display (26) equals $E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ by setting $k=u$ in Lemma C.2. This completes the proof.
Lemma C.4
@@
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1\right\}.$ (27)
Proof 4 (Proof of Lemma C.4)
We prove Lemma C.4 by showing that (27) holds when conditioning both sides on $C_{tu}$ for each $1\leq u\leq\Delta$. Lemma C.4 then follows because $\{C_{tu}:1\leq u\leq\Delta\}$ forms a partition of the whole sample space.
The left hand side of (27) conditional on $C_{t\Delta}$ is
@@
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\}$ (28)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{A}_{t+1:t+\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\}$ (29)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\},$ (30)
which is the right hand side of (27) conditional on $C_{t\Delta}$. Here, (28) follows from the definition of $C_{t\Delta}$; (29) follows from $\prod_{j=t+1}^{t+\Delta-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})=\prod_{j=t+1}^{t+\Delta-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{A}_{t+1:t+\Delta-1})$, a consequence of Assumption 1; and (30) follows from Assumption 1.
The left hand side of (27) conditional on $C_{tu}$ for some $11\leq u\leq\Delta-$ 1 is
@@
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (31)
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{u-1},\bar{A}_{t+u:t+\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (32)
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t+\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (33)
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (34)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\},$ (35)
which is the right hand side of (27) conditional on $C_{tu}$. Here, (31) follows from the definition of $C_{tu}$; (32) follows from Lemma C.1; (33) follows from $\prod_{j=t+1}^{t+u-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{u-1},\bar{A}_{t+u:t+\Delta-1})=\prod_{j=t+1}^{t+u-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{A}_{t+1:t+\Delta-1})$, a consequence of Assumption 1; (34) follows from Assumption 1; and (35) follows from the definition of $C_{tu}$. This completes the proof.
@@ Appendix D Proof of Theorem 1
Proof 5 (Proof of Lemma D.1)
We have
@@
$=$ $E\left(E\left[I_{t}e^{-A_{t}S_{t}^{T}\beta^{*}}\left\{Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha+A_{t}S_{t}^{T}\beta^{*}}\right\}M_{t}W_{t}\left\{A_{t}-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\mid H_{t}\right]\right)$ (36)
$=$ $E\left(E\left[e^{-A_{t}S_{t}^{T}\beta^{*}}\left\{Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha+A_{t}S_{t}^{T}\beta^{*}}\right\}M_{t}W_{t}\left\{A_{t}-\tilde{p}_{t}\left(S_{t}\right)\right\}\mid H_{t},I_{t}=1\right]I_{t}S_{t}\right)$ (37)
@@
$-E\left(E\bigg{[}\big{\{}Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha}\big{\}}\tilde{p}_{t}\left(S_{t}\right)\prod_{j=t+1}^{t+\Delta-1}\big{[}\frac{1\left(A_{j}=0\right)}{1-p_{j}\left(H_{j}\right)}\big{]}^{\mathbbm{1}(\max\limits_{1\leq s\leq j-t}R_{t+s}=0)}\mid H_{t},I_{t}=1,A_{t}=0\bigg{]}I_{t}\left\{1-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\right)$ (38)
@@
$\left.\left.-E\Bigg{(}\prod_{j=t+1}^{t+\Delta-1}\bigg{[}\frac{\mathbbm{1}\left(A_{j}=0\right)}{1-p_{j}\left(H_{j}\right)}\bigg{]}^{\mathbbm{1}(\max\limits_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid H_{t},I_{t}=1,A_{t}=0\Bigg{)}\right\}I_{t}\tilde{p}_{t}\left(S_{t}\right)\left\{1-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\right]$ (39)
$=$ $0$ (40)
Here, (36) follows from the law of iterated expectation; (37) and (38) follows from expanding the expectation on $I_{t}$ and $A_{t}$ respectively; (39) follows from collecting terms; and (40) follows from the identifiability result (4) and the linear model (5) in the main paper, and the fact that $\beta^{*}$ is the true parameter value. This completes the proof.
@@
Assumption D.1 implies that $(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}\left(\alpha^{\prime},\beta^{\prime}\right)$ by Theorem $5.9$ of [34]. Assumption D.1 (uniqueness of $\beta^{\prime}$) and Lemma D.1 imply that $\beta^{*}=\beta^{\prime}$. Therefore, $(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}\left(\alpha^{\prime},\beta^{*}\right)$, and in particular $\hat{\beta}$ is consistent for $\beta^{*}$.
Because $U_{i}(\alpha,\beta)$ is continuously differentiable and hence Lipschitz continuous, Theorem $5.21$ of [34] implies that $\sqrt{n}\big{\{}(\hat{\alpha},\hat{\beta})-\left(\alpha^{\prime},\beta^{*}\right)\big{\}}$ is asymptotically normal with mean zero and covariance matrix $\Sigma=\big{\{}E[\dot{U}_{i}(\alpha^{\prime},\beta^{*})]\big{\}}^{-1}E\left[U_{i}(\alpha^{\prime},\beta^{*})U_{i}(\alpha^{\prime},\beta^{*})^{T}\right]\big{\{}E[\dot{U}_{i}(\alpha^{\prime},\beta^{*})]\big{\}}^{-1,T}$.
@@
$\left\{\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1}\left\{\frac{1}{n}\sum_{i=1}^{n}U_{i}(\hat{\alpha},\hat{\beta})U_{i}(\hat{\alpha},\hat{\beta})^{T}\right\}\left\{\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1,T}.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\dot{U}_{i}(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}E[\dot{U}_{i}(\alpha^{\prime},\beta^{*})]$ (41)
@@
$\frac{1}{n}\sum_{i=1}^{n}U_{i}(\hat{\alpha},\hat{\beta})U_{i}(\hat{\alpha},\hat{\beta})^{T}\stackrel{{\scriptstyle p}}{{\to}}E\left[U_{i}(\alpha^{\prime},\beta^{*})U_{i}(\alpha^{\prime},\beta^{*})^{T}\right],$ (42)
@@
$=[\mathbbm{P}_{n}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}]+[\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\alpha,\beta)\}].$ (43)
The first term in (43) can be controlled by Glivenko-Cantelli Theorem:
@@
because whichever of $\dot{U}_{i}(\alpha,\beta)$ or $U_{i}(\alpha,\beta)U_{i}(\alpha,\beta)^{T}$ that we set $f_{i}$ to, it takes value in a Glivenko-Cantelli class as $(\alpha,\beta)$ is finite-dimensional and takes value in a compact set (Assumption D.1). The second term in (43), $\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\alpha,\beta)\}$, is $o_{p}(1)$ by the dominated convergence theorem because of $(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}\left(\alpha^{\prime},\beta^{*}\right)$ and Assumption D.3. Therefore, (43) is $o_{p}(1)$ for when $f_{i}=\dot{U}_{i}$ and when $f_{i}=U_{i}U_{i}^{T}$. Thus (41) and (42) are proven. This completes the proof of Theorem 1.
Appendix E Drivers of Efficiency Gain

### v3

@@ Appendix C Causal Assumptions and Parameter Identification
Therefore, to prove (C.9) it suffices to show that for any $1\leq u\leq\Delta$,
@@
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$ (C.10)
@@
So right hand side of (C.10) equals
$E\left\{\prod_{j=t+1}^{t+u-1}\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$ (C.11)
Display (C.11) equals $E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ by setting $k=u$ in Lemma 4. This completes the proof.
Lemma 6
@@
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1\right\}.$ (C.12)
Proof 4 (Proof of Lemma 6)
We prove Lemma 6 by showing that (C.12) holds when conditioning both sides on $C_{tu}$ for each $1\leq u\leq\Delta$. Lemma 6 then follows because $\{C_{tu}:1\leq u\leq\Delta\}$ forms a partition of the whole sample space.
The left hand side of (C.12) conditional on $C_{t\Delta}$ is
@@
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\}$ (C.13)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{A}_{t+1:t+\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\}$ (C.14)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1,C_{t\Delta}\right\},$ (C.15)
which is the right hand side of (C.12) conditional on $C_{t\Delta}$. Here, (C.13) follows from the definition of $C_{t\Delta}$; (C.14) follows from $\prod_{j=t+1}^{t+\Delta-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{\Delta-1})=\prod_{j=t+1}^{t+\Delta-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{A}_{t+1:t+\Delta-1})$, a consequence of Assumption 1; and (C.15) follows from Assumption 1.
The left hand side of (C.12) conditional on $C_{tu}$ for some $11\leq u\leq\Delta-$ 1 is
@@
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (C.16)
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{u-1},\bar{A}_{t+u:t+\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (C.17)
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}(\bar{A}_{t+\Delta-1})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (C.18)
$=$ $E\left\{\prod_{j=t+1}^{t+u-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (C.19)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\},$ (C.20)
which is the right hand side of (C.12) conditional on $C_{tu}$. Here, (C.16) follows from the definition of $C_{tu}$; (C.17) follows from Lemma 3; (C.18) follows from $\prod_{j=t+1}^{t+u-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0}_{u-1},\bar{A}_{t+u:t+\Delta-1})=\prod_{j=t+1}^{t+u-1}\mathbbm{1}(A_{j}=0)Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{A}_{t+1:t+\Delta-1})$, a consequence of Assumption 1; (C.19) follows from Assumption 1; and (C.20) follows from the definition of $C_{tu}$. This completes the proof.
@@ Appendix D Proof of Theorem 1
Proof 5 (Proof of Lemma 7)
We have {adjustwidth}-3em-3em
@@
$=$ $E\left(E\left[I_{t}e^{-A_{t}S_{t}^{T}\beta^{*}}\left\{Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha+A_{t}S_{t}^{T}\beta^{*}}\right\}M_{t}W_{t}\left\{A_{t}-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\mid H_{t}\right]\right)$ (D.1)
$=$ $E\left(E\left[e^{-A_{t}S_{t}^{T}\beta^{*}}\left\{Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha+A_{t}S_{t}^{T}\beta^{*}}\right\}M_{t}W_{t}\left\{A_{t}-\tilde{p}_{t}\left(S_{t}\right)\right\}\mid H_{t},I_{t}=1\right]I_{t}S_{t}\right)$ (D.2)
@@
$-E\left(E\bigg{[}\big{\{}Y_{t,\Delta}-e^{g\left(H_{t}\right)^{T}\alpha}\big{\}}\tilde{p}_{t}\left(S_{t}\right)\prod_{j=t+1}^{t+\Delta-1}\big{[}\frac{1\left(A_{j}=0\right)}{1-p_{j}\left(H_{j}\right)}\big{]}^{\mathbbm{1}(\max\limits_{1\leq s\leq j-t}R_{t+s}=0)}\mid H_{t},I_{t}=1,A_{t}=0\bigg{]}I_{t}\left\{1-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\right)$ (D.3)
@@
$\left.\left.-E\Bigg{(}\prod_{j=t+1}^{t+\Delta-1}\bigg{[}\frac{\mathbbm{1}\left(A_{j}=0\right)}{1-p_{j}\left(H_{j}\right)}\bigg{]}^{\mathbbm{1}(\max\limits_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid H_{t},I_{t}=1,A_{t}=0\Bigg{)}\right\}I_{t}\tilde{p}_{t}\left(S_{t}\right)\left\{1-\tilde{p}_{t}\left(S_{t}\right)\right\}S_{t}\right]$ (D.4)
$=$ $0$ (D.5)
Here, (D.1) follows from the law of iterated expectation; (D.2) and (D.3) follows from expanding the expectation on $I_{t}$ and $A_{t}$ respectively; (D.4) follows from collecting terms; and (D.5) follows from the identifiability result (4) and the linear model (5) in the main paper, and the fact that $\beta^{*}$ is the true parameter value. This completes the proof.
@@
Assumption 5 implies that $(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}\left(\alpha^{\prime},\beta^{\prime}\right)$ by Theorem $5.9$ of Van der Vaart (2000). Assumption 5 (uniqueness of $\beta^{\prime}$) and Lemma 7 imply that $\beta^{*}=\beta^{\prime}$. Therefore, $(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}\left(\alpha^{\prime},\beta^{*}\right)$, and in particular $\hat{\beta}$ is consistent for $\beta^{*}$.
Because $U_{i}(\alpha,\beta)$ is continuously differentiable and hence Lipschitz continuous, Theorem $5.21$ of Van der Vaart (2000) implies that $\sqrt{n}\big{\{}(\hat{\alpha},\hat{\beta})-\left(\alpha^{\prime},\beta^{*}\right)\big{\}}$ is asymptotically normal with mean zero and covariance matrix $\Sigma=\big{\{}E[\partial_{\beta}U_{i}(\alpha^{\prime},\beta^{*})]\big{\}}^{-1}E\left[U_{i}(\alpha^{\prime},\beta^{*})U_{i}(\alpha^{\prime},\beta^{*})^{T}\right]\big{\{}E[\partial_{\beta}U_{i}(\alpha^{\prime},\beta^{*})]\big{\}}^{-1,T}$.
@@
$\left\{\frac{1}{n}\sum_{i=1}^{n}\partial_{\beta}U_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1}\left\{\frac{1}{n}\sum_{i=1}^{n}U_{i}(\hat{\alpha},\hat{\beta})U_{i}(\hat{\alpha},\hat{\beta})^{T}\right\}\left\{\frac{1}{n}\sum_{i=1}^{n}\partial_{\beta}U_{i}(\hat{\alpha},\hat{\beta})\right\}^{-1,T}.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\partial_{\beta}U_{i}(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}E[\partial_{\beta}U_{i}(\alpha^{\prime},\beta^{*})]$ (D.6)
@@
$\frac{1}{n}\sum_{i=1}^{n}U_{i}(\hat{\alpha},\hat{\beta})U_{i}(\hat{\alpha},\hat{\beta})^{T}\stackrel{{\scriptstyle p}}{{\to}}E\left[U_{i}(\alpha^{\prime},\beta^{*})U_{i}(\alpha^{\prime},\beta^{*})^{T}\right],$ (D.7)
@@
$=[\mathbbm{P}_{n}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}]+[\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\alpha,\beta)\}].$ (D.8)
The first term in (D.8) can be controlled by Glivenko-Cantelli Theorem:
@@
because whichever of $\partial_{\beta}U_{i}(\alpha,\beta)$ or $U_{i}(\alpha,\beta)U_{i}(\alpha,\beta)^{T}$ that we set $f_{i}$ to, it takes value in a Glivenko-Cantelli class as $(\alpha,\beta)$ is finite-dimensional and takes value in a compact set (Assumption 5). The second term in (D.8), $\mathbbm{P}\{f_{i}(\hat{\alpha},\hat{\beta})\}-\mathbbm{P}\{f_{i}(\alpha,\beta)\}$, is $o_{p}(1)$ by the dominated convergence theorem because of $(\hat{\alpha},\hat{\beta})\stackrel{{\scriptstyle p}}{{\to}}\left(\alpha^{\prime},\beta^{*}\right)$ and Assumption 7. Therefore, (D.8) is $o_{p}(1)$ for when $f_{i}=\partial_{\beta}U_{i}$ and when $f_{i}=U_{i}U_{i}^{T}$. Thus (D.6) and (D.7) are proven. This completes the proof of Theorem 1.
Appendix E More Details on the Efficiency Gain by pd-EMEE
E.1 Elaboration on the Intuition Behind the Efficiency Gain of pd-EMEE
This subsection elaborates the discussion of the efficiency gain of pd-EMEE over EMEE in Section 3. Let us examine how the EMEE weighs the $(i,t)$-decision point by $W_{it}^{\prime}$ and later contrast with our estimator. Case 1: suppose for some $t+1\leq j\leq t+\Delta-1$, $A_{ij}=1$ (Table 4, Examples 1a and 1b). Because this is different from the treatment trajectory for the excursion, the $(i,t)$-decision point is assigned weight $W_{it}^{\prime}=0$ by EMEE. Case 2: suppose $A_{i,t+1}=\cdots=A_{i,t+\Delta-1}=0$ (Table 4, Examples 2a and 2b). In this case, $W_{it}^{\prime}>1$ and the $(i,t)$-decision point is upweighted to account for itself and similar decision points that fall into Case 1 and thus have weight 0.
The key innovation in our proposed estimator is $W_{it}$, which improves efficiency over EMEE. We call $W_{it}$ the per-decision inverse probability weight, and $\hat{\beta}$ the pd-EMEE method, because the inclusion of each of the $(\Delta-1)$ IPW terms is determined by each decision point: for each $j$ from $t+1$ to $t+\Delta-1$, if $\max_{1\leq s\leq j-t}R_{i,t+s}=0$ then $\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}\left(H_{ij}\right)}$ is included in $W_{it}$; otherwise it is not included. Therefore, $W_{it}$ includes in its product the IPW terms from $j=t+1$ to right before the first occurrence of some $R_{j^{*}}=1$. By not including the IPW terms after this $R_{j^{*}}=1$, the per-decision IPW does not account for the treatment assignments after $j^{*}$. This is appropriate because the maximum property (3) implies that the treatment assignments after $j^{*}$ don’t affect the value of $Y_{t,\Delta}$ (which has to equal 1). To see how our estimator improves efficiency over the EMEE, let us now examine how our estimator weighs the $(i,t)$-decision point by $W_{it}$. For Case 1 described above, if some $R_{j}=1$ occurs before the first $A_{k}=1$ (Table 4, Example 1b), the $(i,t)$-decision point will have a positive weight. For Case 2 described above, if the first $R_{j^{*}}=1$ occurs for $j^{*}<t+\Delta-1$ (Table 4, Example 2b), then $W_{it}$ will include fewer terms than $W_{it}^{\prime}$ and thus be smaller, because $W_{it}=\prod_{j=t+1}^{j^{*}}\frac{1}{1-p_{j}\left(H_{ij}\right)}<\prod_{j=t+1}^{t+\Delta-1}\frac{1}{1-p_{j}\left(H_{ij}\right)}=W_{it}^{\prime}$. When the proximal outcome window is shorter than the distance between two decision points so that $\Delta=1$, $\hat{\beta}$ is equivalent to EMEE.
To summarize, by capitalizing on the maximum property (3), the per-decision IP weight $W_{it}$ makes use of some decision points that would be otherwise discarded because $W_{it}^{\prime}=0$ (such as Example 1b) and reduces the weight of some other decision points (such as Example 2b), and this leads to the efficiency improvement. The efficiency improvement of $\hat{\beta}$ over EMEE is larger when a larger proportion of the decision points belong to the above two types, which happens when the proximal outcome window length $\Delta$ is large, when the randomization probability is large, or when the baseline success probability (i.e., the success probability of the proximal outcome under no treatment) is large.
$R_{t}$ $A_{t+1}$ $R_{t+1}$ $A_{t+2}$ $R_{t+2}$ $Y_{t,\Delta}$ $W_{t}$ $W_{t}^{\prime}$
Example 1a 0 0 0 1 0 0 0 0
Example 1b 0 0 1 1 0 1 $\frac{1}{1-p_{t+1}\left(H_{t+1}\right)}$ 0
E.2 A Theoretical Analysis of the Relative Efficiency